The mean price of computers at an electronics store is $850, with a standard deviation of $90. A random sample of 21 televisions is drawn from this population. What is the probability that the mean price of the sample is less than $830?
Answer by stanbon(75887)

The mean price of computers at an electronics store is $850, with a standard deviation of $90.
----------------------------------------
A random sample of 21 televisions is drawn from this population. What is the probability that the mean price of the sample is less than $830?
-----------------------------------------
mean of the distribution of sample means = 850
std of the distribution of sample means = 90/sqrt(21)
---------------
t(830) = (830-850)/[0.90/sqrt(21)] = -1.0184
---
P(x< 830) = P(t< -1.0184 when df=20) = 0.1603
=================================================
Cheers,
Stan H.